OpenStudy (angel_kitty12):

Please please help ASAP I thought I knew this but..nope. (2^2/3•3^2/3)^-6 Also I want to check my answer for this problem 8^-1/2 Thanks so much.

OpenStudy (campbell_st):

so the problem is really \[(\frac{2^2}{3} \times \frac{3^2}{3})^{-6}\] which becomes \[(\frac{2^2 \times 3^2}{3 \times 3})^{-6} \] when dividing the same base subtract the powers \[(2^2 \times 3^{2 - 1 - 1})^{-6} = (2^2 \times 3^0)^{-6}\] so you now need to know about power of zero being 1 so the problem is really \[(2^2)^{-6}\] to simplify this, use power of a power, and multiply the powers. then you may need to recognise the negative power indicates a fraction. hope it helps